OpenStudy (anonymous):

Algebra 2A help please! Find the pH level, to the nearest tenth, of a liquid with [H+] about 3.2x10^-2 a. -2.5 b. 2.5 c. -1.5 d. 1.5 Please and thank you.

OpenStudy (anonymous):

@satellite73

Directrix (directrix):

pH = - log [H+] pH = - log ( 3.2 * 10^(-2) ) pH = - log 3.2 - log 10 ^ (-2) pH = -.505 - (-2) pH = -.51 + 2 pH = 1.49 or 1.5 @fatkat222

OpenStudy (denisaboichuk):

b